# Assessing Inf luential Observations in Pain Prediction using fMRI Data

### v1

@@ 3.1.1 High-dimensional Inf luence Measure (HIM)
Based on the linear regression model (1), Zhao et al. (2013) proposed a non-parametric, model-free and theoretically justified diagnostic measure capable of indirectly assessing the influence of a single observation on variable selection for high-dimensional data. Known as the high-dimensional influence measure (HIM), this metric is formulated on the leave-one-out basis and is designed to quantify the contribution of each individual observation to the marginal correlation between the response variable and all the predictors. An observation with excessively large marginal correlation is deemed influential. To compute the HIM, we first calibrate the influence of the $i^{\text{th}}$ observation on the marginal correlation between the response variable Y and $j^{\text{th}}$ predictor $x_{j}$ by calculating its sample estimate $\widehat{\rho}_{j}^{(i)}=\sum_{k=1,k\neq i}^{n}\big{\{}(\text{X}_{kj}-\widehat{\mu}_{xj}^{(i)})(\text{Y}_{k}-\widehat{\mu}_{y}^{(i)})\big{\}}/\big{\{}(n-1)\widehat{\sigma}_{xj}^{(i)}\widehat{\sigma}_{y}^{(i)}\big{\}}$, $j=1,...,p$ and $i=1,...,n$, where $\widehat{\mu}_{xj}^{(i)}=\sum_{k=1,k\neq i}^{n}\text{X}_{kj}/n$, $\widehat{\mu}_{y}^{(i)}=\sum_{k=1,k\neq i}^{n}\text{Y}_{k}/n$, $\big{(}\widehat{\sigma}_{xj}^{(i)}\big{)}^{2}=\sum_{k=1,k\neq i}^{n}(\text{X}_{kj}-\widehat{\mu}_{xj}^{(i)})^{2}/(n-1)$, and $\big{(}\widehat{\sigma}_{y}^{(i)}\big{)}^{2}=\sum_{k=1,k\neq i}^{n}(\text{Y}_{k}-\widehat{\mu}_{y}^{(i)})^{2}/(n-1)$. Then, upon aggregating and averaging over all $p$ predictors, the HIM for the $i^{\text{th}}$ observation is given by $D_{i}=\sum_{j=1}^{p}\big{(}\widehat{\rho}_{j}-\widehat{\rho}_{j}^{(i)}\big{)}^{2}/p,\ i=1,...,n$, where $\widehat{\rho}_{j}^{(i)}$ is given above and $\widehat{\rho}_{j}$ is its counterpart computed based on the full dataset. Under certain conditions (ibid., Conditions C.1 - C.4, Section 2.3), Zhao et al. (2013) establish that in the absence of influential points, $n^{2}D_{i}\rightarrow\chi^{2}(1)\ \text{as}\ \text{min}(n,p)\rightarrow\infty,\ i=1,...,n$, where $\chi^{2}(1)$ is the chi-squares distribution with one degree of freedom. Thus, the scaled percentile of $\chi^{2}(1)$ can be used as a threshold to pinpoint influential observations. The authors then demonstrate the potential to extend their framework to GLM, yet remarking that the above asymptotic distributional property may not hold. The effectiveness of this method in assessing influence on variable selection is further validated via simulated studies.
@@
Despite advantages including computational simplicity and sound theoretical justification, there are limitations of the HIM approach. First, since it is primarily proposed under the linear regression model, its effectiveness when the response variable is nonlinearly correlated with the predictors is ambiguous. Second, when $p$ is close to $n$, spurious correlation may exist between the response variable and predictors leading to unreliable variable selection, an undesirable phenomenon known as the Freedman’s paradox (Freedman, 1983) which persists in high-dimensional settings (Fan et al., 2014). Thus, adopting marginal correlation to evaluate influence on variable selection may be inappropriate. Third, the HIM approach is independent of model selector, thus providing no insights of the sensitivity of a model selector in spite of the fact that it is not intended to perform such function.
3.1.2 Difference in Model Selection Measured by LASSO (DF(LASSO))
@@
In comparison with the indirect approach by Zhao et al. (2013), under (1), Rajaratnam et al. (2019) formulated the so-called difference in model selection (DF(LASSO)) measure to directly gauge the influence of a single point on the submodel selected by LASSO. Similar to the HIM, the DF(LASSO) also hinges on the leave-one-out scheme and is defined by
$\iota_{i}=\frac{\delta_{i}-\mathbb{E}(\delta_{i})}{\sqrt{\mathrm{Var}(\delta_{i})}},\ i=1,...n,$ (2)
where
@@
Here, $\mathbbm{1}(\cdot)$ is the indicator function, $\widehat{\beta}_{\text{LASSO},j}$ and $\widehat{\beta}^{(i)}_{\text{LASSO},j}$ respectively denote the $j^{\text{th}}$ component of the LASSO estimates obtained from the full dataset and reduced dataset with $i^{\text{th}}$ observation removed. Then, assuming $\sum_{j=1}^{p}\mathbbm{1}(\beta_{j}=0)\rightarrow\infty$, the authors show that
$\frac{\Delta-\mathbb{E}(\Delta)}{\sqrt{\mathrm{Var}(\Delta)}}\rightarrow N(0,1)\ \text{as}\ \text{min}(n,p)\rightarrow\infty,$ (4)
where $\Delta=\sum_{j=1}^{p}\big{\lvert}\mathbbm{1}(\widehat{\beta}_{\text{LASSO},j}=0)-\mathbbm{1}(\beta_{j}=0)\big{\rvert}$ and $N(0,1)$ is a standard Gaussian.
In practice, the sample mean and variance of the $\delta_{i}$’s in (3) are suggested to estimate $\mathbb{E}(\delta_{i})$ and $\mathrm{Var}(\delta_{i})$ in (2). For diagnosis, an observation $i$ with $\lvert\iota_{i}\rvert\geq 2$ is considered influential based on the asymptotic Gaussian approximation in (4). From the formulation of the $\delta_{i}$’s, we see that they play central roles in the assessment as they directly quantify the fluctuation in model selection. Moreover, the comparative magnitude of each $\delta_{i}$ in relation to others, rather than its absolute scale, is more important for decision making. In other words, $\delta_{i}$’s shall be viewed from a collective angle to ascertain the nature of any individual observation. Indeed, this interpretation underpins an important perspective that every individual point is influential on variable selection, yet some may be more influential than the others.
@@
Apart from the above characteristics, there are limitations of the DF(LASSO) framework. First, in establishing (4), the authors recommend using the sample mean and variance of $\delta_{[n]}$ to estimate $\mathbb{E}(\delta_{i})$ and $\mathrm{Var}(\delta_{i})$, which is not theoretically substantiated. Second, the asymptotic normality in (4) is specifically derived for the LASSO estimator. The applicability of this theoretical framework to other model selectors is yet to be established.
@@ 3.1.3 Generalized Difference in Model Selection Measure (GDF)
We now address two limitations associated with the DF(LASSO) approach, namely: (i) the legitimacy of estimating $\mathbb{E}(\delta_{i})$ and $\mathrm{Var}(\delta_{i})$ by the sample mean and variance, and (ii) applicability to extend the framework to other model selectors. To achieve that, we note that $\delta_{[n]}$ is a sequence of dependent random variables and each $\delta_{i}$ is a sum of correlated Bernoulli random variables, where both dependence structures have no tractable forms.
To address the first issue regarding the eligibility of estimating $\mathbb{E}(\delta_{i})$ and $\mathrm{Var}(\delta_{i})$ by the corresponding sample mean and variance, we first focus on the finite-sample property of $\delta_{[n]}$ and bridge the gap between this sequence and the characteristic of the original dataset. With Theorem 1 reported below, we show that $\delta_{[n]}$ is exchangeable and thus identically distributed. The proof of Theorem 1 is given in Section 2 of the Supplementary Materials.
@@
[…]
From Theorem 1, we next provide two relevant results of an exchangeable sequence of random variables in Corollary 1, which are the consistency of the empirical cumulative distribution function (ECDF) and strong law of large numbers (SLLN). The combined statements further strengthen the theoretical basis for estimating $\mathbb{E}(\delta_{i})$ and $\mathrm{Var}(\delta_{i})$ by the sample mean and variance. Its proof is given in Section 3 of the Supplementary Materials.
@@
where $\widehat{\beta}_{j}$ and $\widehat{\beta}^{(i)}_{j}$ are the $j^{\text{th}}$ components of a generic sparse regression coefficient estimate obtained based on the full and reduced data with the $i^{\text{th}}$ observation removed, respectively. As such, $\tau_{i}$ in (5) is referred to as the generalized difference in model selection (GDF) measure in the sequel. To understand the distributional properties of $\tau_{i}$, both parametric and non-parametric approaches can provide reliable and theoretically appropriate approximations. In this article, we specifically focus on one such avenue by theoretically justifying the central limit theorem (CLT) for each $\tau_{i}$ via an approach distinctive from Rajaratnam et al. (2019). Such proposition is further strengthened by additional plausibility arguments via an array of unique angles.
To verify the CLT for $\tau_{i}$, we bridge the theoretical hiatus in Rajaratnam et al. (2019) for generic model selectors through the work of Galambos (1973), which bypasses the challenge in handling the intractable dependence among its summands $\xi_{i,j}$. Specifically, it is shown [ibid., Theorem 1] that for an arbitrary sequence of events, there always exists a sequence of exchangeable events such that the respective sums are equally distributed in finite sample. This implies that $\tau_{i}$ in (5), the sum of dependent indicators $\xi_{i,j},j=1,...,p$, is identically distributed as the sum of a sequence of exchangeable Bernoulli random variables. Toward that end, the regular (Klass & Teicher, 1987) and empirical (Hahn & Zhang, 1998) CLT for an exchangeable sequence of random variables collectively facilitate the derivation of the following assertion, where its proof is available in Section 4 of the Supplementary Materials:
@@
Let $\tau_{i}$ be defined as in (5). If there exists sequences $a_{p}\in\mathbb{R}$, $0<b_{p}\rightarrow\infty$ and $\epsilon_{p}\downarrow 0$ such that as $p\rightarrow\infty$, the following conditions hold: (C.1) $p\cdot\mathbb{P}^{\omega}(\lvert\Xi_{1}\rvert>\epsilon_{p}b_{p})\rightarrow 0$, where $\mathbb{P}^{\omega}$ is a regular conditional distribution for $\xi_{i,j}$ given $\mathcal{F}_{\infty}$, the tail $\sigma$-algebra generated by $\xi_{i,j}$, such that for each $\omega\in\Omega$, the so-called mixands $\Xi_{k}$, $k=1,...,p$, are i.i.d.; (C.2) $b_{p}/\sqrt{p}$ is slowly varying with $v_{p}(\omega)/b_{p}\rightarrow 1$ and $(a_{p}(\omega)-a_{p})/b_{p}\rightarrow 0$, where $a_{p}(\omega)=p\cdot\mathbb{E}\Xi_{1}\mathbbm{1}(\lvert\Xi_{1}\rvert\leq\epsilon_{p}b_{p})$ and $v_{p}^{2}(\omega)=p\cdot\big{(}\mathbb{E}(\Xi_{1}^{2}\mathbbm{1}(\lvert\Xi_{1}\rvert\leq\epsilon_{p}b_{p}))-\big{(}\mathbb{E}\Xi_{1}\mathbbm{1}(\lvert\Xi_{1}\rvert\leq\epsilon_{p}b_{p})\big{)}^{2}\big{)}$, then,
[…]
where $\bar{\xi}$ and $\sigma^{2}_{\xi}$ are the sample mean and variance of the binary sequence $\xi_{i,j}$, $j=1,...,p$.
In practice, since $\tau_{[n]}$ itself is exchangeable by Theorem 1 and with further justification from Corollary 1, $\bar{\xi}$ and $V(\xi)$ can be reasonably supplanted by the sample mean ($\bar{\tau}$) and variance ($\sigma^{2}_{\tau}$) of $\tau_{[n]}$ for large $n$. From (6), at a given nominal level $\alpha\in(0,1)$, the $100\cdot(1-\alpha/2)^{\text{th}}$ percentile of $N(0,1)$ is a valid threshold such that an observation with an absolute magnitude of the standardized $\tau_{i}$ exceeding that cut-off is deemed influential. On the other hand, to test the two-sided hypothesis that the $i^{\text{th}}$ point is not influential, the $p$-value is $2\cdot\mathbb{P}(N(0,1)>(\tau_{i}-\bar{\tau})/\sigma_{\tau})$. Such scheme to identify a single influential observation via the GDF measure is summarized in Algorithm 1.
Algorithm 1 Single Influential Point Detection via the GDF Measure (5)
1:Compute the generic, sparse regression coefficient estimate $\widehat{\boldsymbol{\beta}}$ from the full dataset.
2:Compute $\widehat{\boldsymbol{\beta}}^{(i)}$, $i=1,...,n$, the regression coefficients based on the leave-one-out scheme.
3:Compute $\tau_{i}$, $i=1,...,n$, the associated sample mean ($\bar{\tau}$) and variance ($\sigma^{2}_{\tau}$). Here, $\tau_{i}=\sum_{j=1}^{p}\lvert\mathbbm{1}(\widehat{\beta}_{j}=0)-\mathbbm{1}(\widehat{\beta}^{(i)}_{j}=0)\rvert$, where $\widehat{\beta}_{j}$ and $\widehat{\beta}^{(i)}_{j}$ are obtained from Steps 1 and 2.
4:Given a nominal level $\alpha\in(0,1)$, the individual $k^{\text{th}}$ observation is influential if $\rvert(\tau_{k}-\bar{\tau})/\sigma_{\tau}|>\gamma_{\alpha}$, where $\gamma_{\alpha}$ is the $100\cdot(1-\alpha/2)^{\text{th}}$ percentile of a standard normal $N(0,1)$. To test hypothesis that $k^{\text{th}}$ point is not influential, the $p$-value is $2\cdot\mathbb{P}(N(0,1)>(\tau_{k}-\bar{\tau})/\sigma_{\tau})$.
To further reinforce (6), by de Finetti’s Theorem (de Finetti, 1931), an exchangeable sequence of random variables can be essentially treated as an i.i.d. sequence given its data generating mechanism. Moreover, Aldous (1977) and references therein further explain that a sequence of arbitrarily dependent random variables, including an exchangeable sequence, contains a subsequence sharing analogous properties of its i.i.d. counterpart. Thus, these discussions offer extra plausibility evidence supporting the SLLN and CLT of $\delta_{[n]}$ and $\tau_{[n]}$.

### v2

@@ 4.1.1 High-dimensional Inf luence Measure (HIM)
Based on the linear regression model (1), [37] proposed a non-parametric, model-free and theoretically justified diagnostic measure capable of indirectly assessing the influence of a single observation on variable selection for high-dimensional data. Known as the high-dimensional influence measure (HIM), this metric is formulated on the leave-one-out basis and is designed to quantify the contribution of each individual observation to the marginal correlation between the response variable and all the predictors. An observation with excessively large marginal correlation is deemed influential. To compute the HIM, we first calibrate the influence of the $i^{\text{th}}$ observation on the marginal correlation between the response variable Y and $j^{\text{th}}$ predictor $x_{j}$ by calculating its sample estimate $\widehat{\rho}_{j}^{(i)}=\sum_{k=1,k\neq i}^{n}\big{\{}(\text{X}_{kj}-\widehat{\mu}_{xj}^{(i)})(\text{Y}_{k}-\widehat{\mu}_{y}^{(i)})\big{\}}/\big{\{}(n-1)\widehat{\sigma}_{xj}^{(i)}\widehat{\sigma}_{y}^{(i)}\big{\}}$, $j=1,...,p$ and $i=1,...,n$, where $\widehat{\mu}_{xj}^{(i)}=\sum_{k=1,k\neq i}^{n}\text{X}_{kj}/n$, $\widehat{\mu}_{y}^{(i)}=\sum_{k=1,k\neq i}^{n}\text{Y}_{k}/n$, $\big{(}\widehat{\sigma}_{xj}^{(i)}\big{)}^{2}=\sum_{k=1,k\neq i}^{n}(\text{X}_{kj}-\widehat{\mu}_{xj}^{(i)})^{2}/(n-1)$, and $\big{(}\widehat{\sigma}_{y}^{(i)}\big{)}^{2}=\sum_{k=1,k\neq i}^{n}(\text{Y}_{k}-\widehat{\mu}_{y}^{(i)})^{2}/(n-1)$. Then, upon aggregating and averaging over all $p$ predictors, the HIM for the $i^{\text{th}}$ observation is given by $D_{i}=\sum_{j=1}^{p}\big{(}\widehat{\rho}_{j}-\widehat{\rho}_{j}^{(i)}\big{)}^{2}/p,\ i=1,...,n$, where $\widehat{\rho}_{j}^{(i)}$ is given above and $\widehat{\rho}_{j}$ is its counterpart computed based on the full dataset. Under certain conditions (ibid., Conditions C.1 - C.4, Section 2.3), [37] establish that in the absence of influential points, $n^{2}D_{i}\rightarrow\chi^{2}(1)\ \text{as}\ \text{min}(n,p)\rightarrow\infty,\ i=1,...,n$, where $\chi^{2}(1)$ is the chi-squares distribution with one degree of freedom. Thus, the scaled percentile of $\chi^{2}(1)$ can be used as a threshold to pinpoint influential observations. The authors then demonstrate the potential to extend their framework to GLM, yet remarking that the above asymptotic distributional property may not hold. The effectiveness of this method in assessing influence on variable selection is further validated via simulated studies.
@@
Despite advantages including computational simplicity and sound theoretical justification, there are limitations of the HIM approach. First, since it is primarily proposed under the linear regression model, its effectiveness when the response variable is nonlinearly correlated with the predictors is ambiguous. Second, when $p$ is close to $n$, spurious correlation may exist between the response variable and predictors leading to unreliable variable selection, an undesirable phenomenon known as the Freedman’s paradox [12] which persists in high-dimensional settings [11]. Thus, adopting marginal correlation to evaluate influence on variable selection may be inappropriate. Third, the HIM approach is independent of model selector, thus providing no insights of the sensitivity of a model selector in spite of the fact that it is not intended to perform such function.
4.1.2 Difference in Model Selection via the LASSO (DF(LASSO))
@@
In comparison with [37], under (1), [28] formulated the so-called difference in model selection (DF(LASSO)) measure to directly gauge the influence of a single point on the submodel selected by the LASSO. Similar to the HIM, the DF(LASSO) is also based on the leave-one-out scheme and is defined by $\iota_{i}=(\delta_{i}-\mathbb{E}(\delta_{i}))/\sqrt{\mathrm{Var}(\delta_{i})}$, $i=1,...,n$, where
@@
Here, $\mathbbm{1}(\cdot)$ is the indicator function, $\widehat{\beta}_{\text{LASSO},j}$ and $\widehat{\beta}^{(i)}_{\text{LASSO},j}$ respectively denote the $j^{\text{th}}$ component of the LASSO estimates obtained from the full dataset and reduced dataset with $i^{\text{th}}$ observation removed. Then, assuming $\sum_{j=1}^{p}\mathbbm{1}(\beta_{j}=0)\rightarrow\infty$, the authors verify that $(\Delta-\mathbb{E}(\Delta))/\sqrt{\mathrm{Var}(\Delta)}\rightarrow N(0,1)\ \text{as}\ \text{min}(n,p)\rightarrow\infty$, where $\Delta=\sum_{j=1}^{p}\big{\lvert}\mathbbm{1}(\widehat{\beta}_{\text{LASSO},j}=0)-\mathbbm{1}(\beta_{j}=0)\big{\rvert}$ and $N(0,1)$ is a standard Gaussian.
In practice, the sample mean and variance of the $\delta_{i}$’s in (2) are suggested to estimate $\mathbb{E}(\delta_{i})$ and $\mathrm{Var}(\delta_{i})$ in the DF(LASSO) measure. For diagnosis, an observation $i$ with $\lvert\iota_{i}\rvert\geq 2$ is considered influential by the asymptotic Gaussian approximation. From the formulation of the $\delta_{i}$’s, we see that they play central roles in the assessment as they directly quantify the fluctuation in model selection. Moreover, the comparative magnitude of each $\delta_{i}$ in relation to others, rather than its absolute scale, is more important for decision making. In other words, $\delta_{i}$’s shall be viewed from a collective angle to ascertain the nature of any individual observation. Indeed, this interpretation underpins an important perspective that every individual point is influential on variable selection, yet some may be more influential than the others.
@@
Apart from the above characteristics, there are limitations of the DF(LASSO) framework. First, in establishing the asymptotic normality, the authors recommend using the sample mean and variance of $\delta_{[n]}$ to estimate $\mathbb{E}(\delta_{i})$ and $\mathrm{Var}(\delta_{i})$, which is not theoretically substantiated. Second, such asymptotic normality is specifically derived for the LASSO estimator. The applicability of this theoretical framework to other model selectors is yet to be established.
@@ 4.1.3 Generalized Difference in Model Selection Measure (GDF)
We now develop a universally feasible diagnostic measure by addressing two limitations of the DF(LASSO) method, namely: (i) the legitimacy of estimating $\mathbb{E}(\delta_{i})$ and $\mathrm{Var}(\delta_{i})$ by the sample mean and variance, and (ii) applicability to extend the framework to other model selectors. To achieve that, we note that $\delta_{[n]}$ is a sequence of dependent random variables and each $\delta_{i}$ is a sum of correlated Bernoulli random variables, where both dependence have no tractable forms.
To address the first issue regarding the eligibility of estimating $\mathbb{E}(\delta_{i})$ and $\mathrm{Var}(\delta_{i})$ by the corresponding sample mean and variance, we first focus on the finite-sample property of $\delta_{[n]}$ and bridge the gap between this sequence and the characteristic of the original dataset. With Theorem 1 reported below, we show that $\delta_{[n]}$ is exchangeable and thus identically distributed. The proof of Theorem 1 is given in Section 2 of the Supplementary Materials Part I.
@@
Suppose that the dataset $\textbf{Z}=(\text{Z}_{1},...,\text{Z}_{n})^{\top}$, where $\text{Z}_{i}=(\text{Y}_{i},\textbf{X}_{i}^{\top})^{\top}\in\mathbb{R}^{p+1},\ i=1,...n$, has i.i.d. rows. Then, the sequence of non-negative, discrete diagnostic measures $\delta_{[n]}$ defined in (2) is exchangeable and thus marginally identically distributed.
From Theorem 1, we next provide two relevant results of an exchangeable sequence of random variables in Corollary 1, which are the consistency of the empirical cumulative distribution function (ECDF) and strong law of large numbers (SLLN). The combined statements further strengthen the theoretical basis for estimating $\mathbb{E}(\delta_{i})$ and $\mathrm{Var}(\delta_{i})$ by the sample mean and variance. Its proof is given in Section 3 of the Supplementary Materials Part I.
@@
where $\widehat{\beta}_{j}$ and $\widehat{\beta}^{(i)}_{j}$ are the $j^{\text{th}}$ components of a generic sparse regression coefficient estimate obtained based on the full and reduced data with the $i^{\text{th}}$ observation removed, respectively. As such, $\tau_{i}$ in (3) is referred to as the generalized difference in model selection (GDF) measure in the sequel. To understand the distributional properties of $\tau_{i}$, both parametric and non-parametric approaches can provide reliable and theoretically appropriate approximations. In this article, we specifically focus on one such avenue by theoretically justifying the central limit theorem (CLT) for each $\tau_{i}$ via an approach distinctive from [28].
To verify the CLT for $\tau_{i}$, we bridge the theoretical hiatus in [28] for generic model selectors through the work of [13], which bypasses the challenge in handling the intractable dependence among its summands $\xi_{i,j}$. Specifically, it is shown [ibid., Theorem 1] that for an arbitrary sequence of events, there always exists a sequence of exchangeable events such that the respective sums are equally distributed in finite sample. This implies that $\tau_{i}$ in (3), the sum of dependent indicators $\xi_{i,j},j=1,...,p$, is identically distributed as the sum of a sequence of exchangeable Bernoulli random variables. Toward that end, the regular [19] and empirical [15] CLT for an exchangeable sequence of random variables collectively facilitate the derivation of the following assertion, where its proof is available in Section 4 of the Supplementary Materials Part I:
@@
Let $\tau_{i}$ be the GDF measure defined as in (3). Suppose that the conditions (A.1) and (A.2) (see Section 4 of the Supplementary Materials Part I) hold. Then,
$\frac{\tau_{i}-\bar{\tau}}{\sigma_{\tau}}\rightarrow N(0,1)\ \text{as}\ \text{min}(n,p)\rightarrow\infty,$ (4)
where $\bar{\tau}$ and $\sigma^{2}_{\tau}$ are the sample mean and variance of the sequence $\tau_{[n]}$.
While we essentially have summation of Bernoulli variables, we have presented the conditions in rather general form to extend the applicability of the result. Here, condition (A.1) controls the tail behavior of the exchangeable summand of $\tau_{i}$ by requiring that it is not heavy tailed, which is a weak version of the existence of the first moment. Condition (A.2) is parallel to the classical Lindeberg’s condition in establishing the CLT, which requires that the contribution of each individual summand to the variance of $\tau_{i}$ is sufficiently small. In fact, the existence of a uniform bound on the second moment of the summand serves as a sufficient (not necessary) condition for (A.1) and (A.2). In our case, each individual summand is Bernoulli so that both conditions are fulfilled. From (4), at a given nominal level $\alpha\in(0,1)$, the $100\cdot(1-\alpha/2)^{\text{th}}$ percentile of $N(0,1)$ is a valid threshold such that a point with an absolute scale of the standardized $\tau_{i}$ exceeding that cut-off is deemed influential. Moreover, to test the two-sided hypothesis that the $i^{\text{th}}$ point is clean, the $p$-value is $2\cdot\mathbb{P}(N(0,1)>(\tau_{i}-\bar{\tau})/\sigma_{\tau})$. A framework for detecting a single influential point via the GDF metric is given in Algorithm GDF in Section 5 of the Supplementary Materials Part I.